Five wrenches, arraged in order of size, have teh following measures inches 3/8, 7/16, 1/2, x, and 5/8.
If the wrenches increase in size by the same amount, what is the measure in inches of x?

Answer by jorel1380(3719)

If they're increasing in size;
7/16-3/8=7/16-6/16=1/16
The next one after 1/2 is 8/16+1/16 or 9/16..